AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 01:28:13 2025
Author
*
:
Subject
*
:
<p>13:37 Run started - no beam in the ring (One of the ring magnets was still on RED when it should have been set to green by HKR)</p> <p>14:28 Beam ON - still observe beam halo on DSSD0 (Top left)</p> <p> YR09 Vertical and horizontal offsets set to -9.0 mm</p> <p>14:36 Target ON, Beam ON</p> <p> We can see the beam halo on the detectors again in the same position as previous tests. The energy histograms for the detectors show only low energy events, except bottom left detector where we see a peak around 20 MeV. With the target on we see additional events but we only see events at ~20MeV for the bottom detector still in the bottom left detector. We are moving the detectors IN to try and see events from scattering of the beam on the target. We may not be hitting the target with the beam.</p> <p> </p> <p>15.29 We observe data in DSSD # 0, 2, 3 (top left, bottom left, bottom right) which is consistent with a line beam halo on the detector. The top right detector is positioned higher by ~10 mm than the top left detector so it is not unphysical to not see any events in this detector from the halo.</p> <p> The energy of events in each detector is not consistent however. In DSSD #0 and 3 we see mainly low energy events (some small number of 20 MeV events is seen) but in DSSD #2 we see mainly higher energy (20 MeV) events. See attached energy histograms, xy and Ex vs x plots (attachments 1-3).</p> <p> This is consistent with the data from channels seen in MIDAS for aida04 (DSSD #0) and aida12 (DSSD#2).</p> <p> Switched FEE64 range bit frpom 20MeV FSR to 1GeV FSR to see if any events are over the 20 MeV range of the low energy channel . Slow comparator from 0x64 to 0x5. We see no events just above 20 MeV in aida04 but do in aida12. See attachments 4 & 5</p> <p> No data ( > 20MeV nominal ) out of range in DSSSD #1</p> <p> FEE range set back to FSR to 20 MeV and slow comparator set back to 0x64</p> <p>15:57 Beam on target. Left arm fully out (pot 18.8 mm), right arm fully in (pot 83.69 mm). Beam injection ~1e6, target density 2e11. Cycle time between 4-13 s.</p> <p> Leaving for more stats</p> <p> </p> <p> </p> <p>17.30 Routine system checks</p> <p> DAQ continues file G22-88/R16_14</p> <p> DSSSD bias and leakage current OK - attachments 8 & 28</p> <p> DSSSD #2 bias -100V, all other DSSSDs -150V</p> <p> ambient temperature +17.0 deg C</p> <p> </p> <p> FEE64 temperatures OK - attachment 9</p> <p> All system wide checks OK</p> <p> WR timestamp OK - attachment 10</p> <p> aida01 & aida03 ASIC settings - attachments 11-12</p> <p> AIDA ADC, PAUSE, DISC, correlation - attachments 13-16</p> <p> ADC data item rates OK</p> <p> per FEE64 Rate spectra - attachments 17-18</p> <p> note change in rates for aida12 & aida13 - presumably due to beam cycle?</p> <p> per FEE64 1.8.L, 1.8.H and 1.8.W 20us FSR spectra - attachments 19-24</p> <p> Merger, Merger stats and TapeServer - attachments 25-27</p> <p>18.10 ML CRYRING vacuum - attachment 29</p> <p>FEE spectrum browser statistics attachment 30 - not many missing channels - aida04 asic 2 perhaps?</p> <p>18:40 End of run. Target x position moved out by 3000. Faraday cup to the ring closed. Detectors moved fully OUT.</p> <p>Run 17 started - alpha background</p> <p> </p> <p> </p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown